OpenStudy (anonymous):

Solve. 2x^2 + x = 15. Answer x = 3/2 and x = –5 x = –5 and x = 5 x = 5/2 and x = –3 x = 15 and x = –2 Solve. 2x^2 + x = 15. Answer x = 3/2 and x = –5 x = –5 and x = 5 x = 5/2 and x = –3 x = 15 and x = –2 @Mathematics

OpenStudy (agreene):

C. x=5/2 and -3

OpenStudy (anonymous):

2x^2 + x - 15 = 0, (2x-5)(x+3) = 0, x = 5/2 or x = -3

OpenStudy (anonymous):

x= -3, 5/2

OpenStudy (anonymous):

Thank you everyone :)